OpenStudy (anonymous):

How many solutions will this system of equations have? 4x-2y=12 3y-6x=-18 a. None b. One c. Two d. Infinite

OpenStudy (anonymous):

4x-2y=12 divide by 2 2x-y=6 -6x+3y=--18 divide by -3 2x-y=6 actually it represents a st. line only and a st .line constitute infinite numbers.